OpenStudy (anonymous):

how do we find the range for root (9-x)

OpenStudy (mr.math):

\[\sqrt{9-x}\ge 0.\]

OpenStudy (mr.math):

For a any polynomial P(x), \(\sqrt{P(x)}\ge 0\).

OpenStudy (mr.math):

For any*

OpenStudy (anonymous):

You can't take the square root of a negative number. So the domain is \[\Large x \in ( - \infty ,9]\]This means that the lowest possible output is 0, when x = 9. So the range is all numbers from 0 to infinity.

OpenStudy (ash2326):

we havy y= root(9-x) let's see the domain of y 9-x>=0 ( expression inside root should be greater than 0) x<=9 so domain is (-infinity, 9] range can be found by puttin the values of x if x=9 y=0 y can't be less than 0 if x=-infinity then y is infinity so range is [0,infinity) or range is>=0

OpenStudy (anonymous):

Hmm OpenStudy is giving me grief... It wasn't showing me your answer before I posted Mr.Math :'(

OpenStudy (anonymous):

I think answer is all non-negative real numbers cause here we have to see that its sqrt operator and range of it is the above said and so the answer is [0,infnty)